OpenStudy (anonymous):

find the all values of x for which f(x)=0 let f(x)=x^-5x+6

mathslover (mathslover):

nhof , is it : f(x) = \(\large x^{-5x +6}\) or f(x) = \(x^2 - 5x + 6\) ?

OpenStudy (precal):

probably the later f(x)=0 means that your f(x) value is zero so let's use substitution and solve for x 0=x^2-5x+6 I would factor the right side and solve each factor. This is a quadratic (ie 2nd power) and you should have 2 solutions.

OpenStudy (anonymous):

x^2-5x+6 (x-2)(x-3) x-2=0 x-3=0 x=2 x=3